OpenStudy (kendricklamar2014):

Given the function rule f(x) = x^2 -5x +1, what is the output of f(-3) A. 7 B. -5 C. 25 D. 5

OpenStudy (kendricklamar2014):

@geekfromthefutur

OpenStudy (ranga):

f(x) = x^2 -5x +1 f(-3) = (-3)^2 - 5(-3) + 1 = ?

OpenStudy (anonymous):

Don't you just have to plug in -3

OpenStudy (kendricklamar2014):

thanks